# Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry

### v3

@@ 3.5. Data Analysis
We analyzed interview transcripts using reflexive thematic analysis (Clarke and Braun, 2021), which provides systematic yet adaptable qualitative data interpretation. Its independence from pre-existing theoretical frameworks suited our exploratory focus on youth-envisioned social media design through FI, enabling effective capture and analysis of diverse participant perspectives.
Since not all participants agreed to have their videos on during interviews, we did not conduct video analysis of gestures that participants sometimes used when describing their ideas. Additionally, drawing was not part of our protocol because we have experienced in prior interviews for other studies that asking participants to draw can create high pressure for teen and youth participants, potentially limiting their imagination. All data therefore consisted of transcripts from verbal interviews.
Analysis began with the first two authors independently coding the same two transcripts line-by-line using descriptive codes closely adhering to data. Authors discussed and resolved coding discrepancies, then individually coded two different transcripts using Atlas.ti (atl, 2024). After four iterations, we reached code saturation. The first author then coded all transcripts using the final code set. All code versions and the final codebook are in supplementary materials.
@@ 3.6. Ethical Considerations
Given J.K. Rowling’s controversial transgender statements (Oth, 2020; Miller, 2024), we included a disclaimer  [^1] The disclaimer read, We acknowledge that J.K. Rowling has made statements regarding transgender individuals that many find harmful and discriminatory. We do not support or endorse these views. Our study focuses on the magical context of the Harry Potter series, and we will strive to create an inclusive and respectful environment for all participants, regardless of gender identity or expression. Thank you for your understanding and for helping us advance this important research. in recruitment materials and surveys, reiterating our position during interviews.
@@ 4. Results
4.1. Intuitive Social Navigation
Youth sought interactions that feel natural, embodied, and socially meaningful, reflecting how people navigate relationships and spaces in everyday life. They imagined SMH not as a collection of buttons, but as a space where presence, movement, and context shape connection, self-expression, and privacy management in intuitive, socially familiar ways.
@@ 4.1.1. Supporting Presence and Connection Through Embodied, Immersive Interaction
Participants described how embodied, immersive qualities like realistic movement, lifelike avatars, and holographic interfaces could transform SMH into a space fostering presence, emotional connection, and meaningful interaction. In contrast, current social media (“Muggle Social Media” or MSM) was described as doing “a good job at spreading information [but not] emotions.” (P16) P01 captured this limitation, explaining that “if you’re just watching things through a screen… technically, there’s like a wall dividing you.”
For others, fulfillment came from embodied physical actions. P04 valued “meandering through different people’s spaces,” finding deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 appreciated how moving physically between rooms created a “feeling of time passing,” explaining “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.” They added that traversal in a “corporal form” would feel “fulfilling,” “calming,” and “reassuring.”
P03 reflected on how simple actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.”
This sense of presence was felt even in seemingly passive actions. As P01 explained,
@@
For P01, 3D interaction supported more embodied engagement and introduced a felt presence that lingered, even without direct interaction.
@@ 4.1.2. Using Spatial Cues and Social Norms to Manage Access and Privacy
Participants proposed using spatial and visual cues modeled after real-world environments to make managing access and privacy feel more intuitive, socially natural, and contextually appropriate. P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
P01 contrasted 2D and 3D settings, noting how boundaries over personal space are clearer: “In 2D, you’re just observing.… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” Therefore, they expected others to recognize and respect spatially cued privacy norms:
@@
While managing visibility across multiple rooms might seem complex, participants believed this approach would feel intuitive and natural. P11 suggested spatial contexts could reduce privacy control efforts, stating “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
4.2. Shared Experiences
Participants emphasized that shared experiences, through active collaboration or simply being together, play a crucial role in forming meaningful relationships and lasting memories.
@@ 4.2.1. Fostering Shared Memories Through Collaborative Activities
Participants described shared activities as meaningful foundations for building friendships. P06 contrasted casual solo games with more immersive, co-experienced ones. While playing a simple emoji game lacked true shared experience since “we don’t actually see each other’s screens,” they described Roblox as fostering lasting memories: “do you remember that time when that kid in the voice chat said this, or when we got scared by that jump scare… it’s not just that we’re in the same space, but that we’re doing something together and experiencing it together.”
P12 described the value of collaborative projects like “creat[ing] a house we love” online, which offered both “a sense of self-achievement” and “teamwork,” contrasting with the passivity of “just like doom scrolling.” For P12, working together on “the same topic, on the same project” felt like “a better way to connect” than “posting” content individually.
@@ 4.2.2. Making Space for Casual Connections Through Low-Pressure, Ambient Social Interaction
Participants saw low-pressure, ambient social interaction, like co-existing in shared spaces or listening to music together, as natural ways to build connections and shared memories. P18 described a “study sesh” where friends could send avatars to “sit and study together” in a virtual study room with Lo-Fi music. These spaces would mirror Discord servers’ 24/7 activity, offering comforting presence even without direct conversation. Even “mindless scrolling” became meaningful when shared, as P18 noted that “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
Music often facilitated these interactions, with P19 envisioning a “jukebox” where friends could “queue up songs,” reflecting on their experience:
@@
They added that music could serve as a “conversation starter” while providing a passive, enjoyable way to connect: “really does bring joy to people because you don’t have to connect through talking to someone about music. You can just listen to the music.”
4.3. Communal Ambience
SMH was envisioned as a platform for intentional, engaging interactions with close friends, making it naturally less prone to social comparison. P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, participants emphasized features fostering organic, intentional, personalized interactions and closer friendships. They described several design features–neighborhood structures, shared gathering spaces, and stronger boundary controls–that could cultivate this communal, low-pressure environment.
@@ 4.3.1. Designing Neighborhood Structures that Reflect Real-World Communities
Participants wanted SMH’s 3D world to mirror real-world communities, using familiar layouts like homes, neighborhoods, and shared areas to foster social interaction and relationship-building. P22 described in detail:
@@
Youth envisioned SMH reflecting real-world relationship dynamics, using physical distances within the landscape to represent social tie strength. Some participants envisioned SMH resembling a college campus or neighborhood, where physical closeness naturally fosters interaction. One participant elaborated:
@@
P01 proposed a tiered layout where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be farther away, requiring a “little hike” to reach. This approach of aligning physical landscapes with emotional closeness would streamline interactions and focus on meaningful relationships. P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.”
@@ 4.3.2. Establishing Third Places for Casual Social Interaction and Serendipitous Connections
Participants imagined SMH neighborhoods including shared third places where people could casually gather, lowering social barriers and creating connection opportunities. P05 described these as “neutral spaces” not tied to specific activities or relationships, where groups could gather for various purposes. They imagined “some places to sit, and some kind of flat table-like surface” supporting casual activities like “doing homework or study, or [playing] some kind of board game.” Such spaces would function as “a third, neutral territory” for organic interaction without managing invitations or switching platforms.
Participants described shared neighborhood spaces like common rooms, courtyards, or clubhouses where residents could casually interact outside personal spaces. P01 described gathering in a “courtyard or game room” at the neighborhood center. Drawing on Club Penguin  and Poptropica , they imagined this shared space as a “clubhouse… open to anybody,” with themed rooms catering to different interests, like a Quidditch field or library, creating opportunities to meet others through shared activities. P09 imagined a “virtual dining room” allowing casual conversation while eating together. Unlike Hogwarts’ segmented house-based accommodation, this dining space was open and inclusive, enabling cross-social group connections.
Others described public third places supporting casual, serendipitous interactions. P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you.”
@@ 4.3.3. Prioritizing Close Relationships and Reducing Extraneous Content
To preserve SMH’s focus on meaningful, stress-free interactions, participants emphasized filtering irrelevant content like celebrity drama or political commentary while prioritizing updates from close friends and trusted communities. P19 criticized how platforms like Instagram have shifted away from their initial focus:
@@
Several participants advocated minimizing celebrity influence, with P01 suggesting isolating them to a “celebrity zone.” P07 captured the preference for close connections: “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” They envisioned SMH being “more designed for people you actually know… like your friends, classmates, or Quidditch team.”
@@ 4.3.4. Developing Socially Grounded and Themed Conversational Spaces
Participants described how themed, socially grounded conversational spaces for protected, meaningful dialogue could make SMH feel more personal and emotionally connected than MSM.
@@
Participants described how conversational space design could make SMH interactions feel safer, more personal, and more inviting than existing platforms. Many valued walking into themed, topic-specific rooms where conversation felt protected and purposeful, rather than broadcasting to broad, undefined audiences. P22, for example, imagined “chat rooms you can walk into,” each centered on themes like “high school stuff,” decorated with artifacts from shared experiences to help spark meaningful conversations. They described these as “protected space[s]…where you can talk…without feeling unsafe or judged,” contrasting them with the more exposed, contextless feel of mainstream social media where “you don’t know everything around it.”
Participants also imagined personalizing these conversational spaces to make them feel more emotionally connected. P07 described filling chatrooms with “photos of you together, inside jokes, or fun memories,” making even casual conversations feel more intimate. P09 expanded on this idea, imagining a “Room of Requirement”-like system (roo, [n. d.]) where users could create custom rooms on demand to match the “purpose” of their conversation. Participants also valued interest-based spaces, similar to Discord (discord page, [n. d.]), where a “common point” (P13) or “good talking point” (P07) could help break the ice and give conversations direction.
@@ 4.3.5. Implementing Comprehensive Boundary Controls to Create Personal Spaces that Feel Safe
To ensure SMH feels safe and personal, participants stressed needing comprehensive boundary controls beyond existing platforms’ limited and inconsistent privacy tools. P22 explained that even with contact syncing off, Instagram still suggested their Finsta (secondary account (Huang and Vitak, 2022)) to people they wanted to avoid, based on mutual follows. They stressed needing stronger, more thorough blocking controls, including blocking “alternate accounts” and fully removing blocked person traces, explaining “I don’t wanna be able to tell that the person I blocked is even around me.” P14 pointed out that even private Instagram accounts expose personal content like Story Highlights to “a random person”, revealing perceived privacy protection gaps (Bojkov, 2019).
Participants wanted stronger control over their visibility, imagining features letting them opt out of being seen or discovered entirely. P03 suggested an “invisibility cloak” for moments when users “don’t want anyone to see you.”
@@
4.4. Nuanced Self-Presentation
Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14). This curated image culture creates “a lot of seriousness” (P23), reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23). P13 shared helping a friend prepare an Instagram post, including selecting “post order,” highlighting “how seriously we take social media and our image online.” In contrast, participants imagined SMH as a space where self-presentation could feel more personal, authentic, and flexible, reflecting their full range without curation or performance pressures.
@@ 4.4.1. Expressing Identity Through Customizable Avatars, Spaces, and Artifacts
Participants consistently described customizable avatars and personal spaces as meaningful self-expression tools, offering far more creativity than MSM’s limited settings. P21 envisioned building their own “mini world” complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school,” sharing these spaces as “showing a little part of yourself.” P03 described wanting to create a “dark academia” Hogwarts-inspired room, filled with “so many bookshelves and… Victorian architecture.”
Customizable “3D avatar[s]” (P11) and profiles were key ways of showing individuality. P22 reflected on early Myspace (mys, [n. d.]), describing how they enjoyed coding profile layouts and changing “the font and really make it reflect me.” Even small details like choosing usernames were ways to “convey to the world and your friends your own identity… which is a positive feeling usually.” P01 noted such expression is often “mainly for you,” not just for public display. P17 added that even if users don’t consciously intend to reveal something personal, their choices still do: “your brain conjured that up, right? So I think that’s part of you.”
Beyond visual customization, participants imagined SMH supporting ambient, indirect expression, offering subtle cues about what someone is doing, feeling, or thinking without formal posts. P01 wanted their avatar to reflect real-time states like current outfit or music, so friends could intuitively notice “‘Oh. that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ P10 imagined friends picking up on interest changes by observing virtual space updates over time, providing “an insight into… what I’m thinking.”
Several participants highlighted the contrast between 2D platforms’ flat, compressed feel and 3D environments’ immersive qualities. P16 noted that VR creates presence feeling more like meeting “real people… compared to profile pictures,” because you can pick up on “mannerisms” and details like “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.”
Participants described how these choices feel more authentic than generic posts because they emerge naturally from personal preferences and interests. As P01 put it, such “creative expression” feels “a lot more personal” because it reflects “that’s not just like what life has given you… a reflection of your outlook on life.” P13 explained that these subtle forms of sharing make social media feel more like “an extension of everyday interaction,” helping people “gain the basic foremost information” about someone without needing to start a direct conversation. P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
Drawing parallels to the physical world, P13 pointed out how personal spaces like bedrooms carry a sense of intimacy and authenticity that translates well to SMH. They explained:
@@
Finally, participants appreciated these ambient and low-pressure forms of expression for their lower social stakes. P16, who described themselves as “fairly private” and not particularly “public” about their preferences, explained that “decorat[ing] a house” or customizing an avatar allows others to “get to know more about my personality… through my style” without feeling like they are performing for attention. They described this kind of expression as “more justified” and less “embarrassing” than posting content just to be seen by others.

### v4

@@ 3.5. Data Analysis
We analyzed interview transcripts using reflexive thematic analysis (BraunClarke-2021-ThematicAnalysisGuide-k), which provides systematic yet adaptable qualitative data interpretation. Its independence from pre-existing theoretical frameworks suited our exploratory focus on youth-envisioned social media design through FI, enabling effective capture and analysis of diverse participant perspectives.
Since not all participants agreed to have their videos on during interviews, we did not conduct gesture analysis. Additionally, drawing was not part of our protocol because we have found in prior interviews that asking participants to draw can create high pressure for teen and youth participants, potentially limiting their imagination. All data therefore consisted of verbal interview transcripts.
Analysis began with the first two authors independently coding the same two transcripts line-by-line using descriptive codes closely adhering to data. Authors discussed and resolved coding discrepancies, then individually coded two different transcripts using Atlas.ti (atlas). After four iterations, we reached code saturation. The first author then coded all transcripts using the final code set. All code versions and the final codebook are in supplementary materials.
@@ 3.6. Ethical Considerations
Given J.K. Rowling’s controversial transgender statements (Other-2020-JKIssues-x; Miller-2024-JKTransantagonism-a), we included a disclaimer in recruitment materials and surveys, reiterating our position during interviews. The disclaimer read: We acknowledge that J.K. Rowling has made statements regarding transgender individuals that many find harmful and discriminatory. We do not support or endorse these views. Our study focuses on the magical context of the Harry Potter series, and we will strive to create an inclusive and respectful environment for all participants, regardless of gender identity or expression. Thank you for your understanding and for helping us advance this important research.
@@ 4. Results
In line with the goals of co-design processes (Sanders2008Co-creationDesignh), our interest lay less in the specifics of functionality and more in uncovering what kinds of interactions and environments participants valued most.
4.1. Theme 1: Intuitive Social Navigation
Youth sought interactions that feel natural, embodied, and socially meaningful—reflecting how people navigate relationships and spaces in everyday life. They imagined SMH not as a collection of buttons, but as a space where presence, movement, and context shape connection, self-expression, and privacy in intuitive, familiar ways.
@@ 4.1.1. Supporting Presence and Connection Through Embodied, Immersive Interaction
Embodied, immersive qualities—realistic movement, lifelike avatars, holographic interfaces—could transform SMH into a space fostering presence and emotional connection. Current social media (“Muggle Social Media” or MSM) was described as doing “a good job at spreading information [but not] emotions.” (P16) P01 captured this limitation: “if you’re just watching things through a screen… technically, there’s like a wall dividing you.”
For others, fulfillment came from embodied physical actions. P04 valued “meandering through different people’s spaces,” finding deliberate exploration heightened their sense of being “more present in somebody’s space.” P09 appreciated how moving physically between rooms created a “feeling of time passing”: “there’s a feeling like you’re actually officially leaving a space as opposed to clicking into a different one.” Traversal in a “corporal form” would feel “fulfilling,” “calming,” and “reassuring.”
P03 reflected that simple actions like “mov[ing] your legs to walk” or “turn[ing] your head” add a “tiny bit of connection,” making it “easier to have connections with people.”
This sense of presence emerged even in seemingly passive actions. As P01 explained,
@@
For P01, 3D interaction introduced a felt presence that lingered, even without direct interaction.
@@ 4.1.2. Using Spatial Cues and Social Norms to Manage Access and Privacy
Participants proposed spatial and visual cues modeled after real-world environments to make access and privacy management feel more intuitive and contextually appropriate. P03 described public areas like “the downstairs area,” “guest bedroom,” or “game room” as spaces where people could “just come in” without permission, while private spaces like their bedroom required “extra consent” as it was a “safe space.” P13 suggested intermediate zones like a “waiting room” for sharing “semi-personal things” before granting further access.
P01 contrasted 2D and 3D settings, noting how boundaries over personal space are clearer: “In 2D, you’re just observing.… But in a 3D setting, even if you don’t leave physical footprints, in a way, they’re still there.” They expected others to recognize and respect spatially cued privacy norms:
@@
While managing visibility across multiple rooms might seem complex, participants believed this approach would feel intuitive. P11 suggested spatial contexts could reduce privacy control efforts: “It would make things easier because you’d be able to easily tell who’s allowed and not allowed in certain areas of the house.”
4.2. Theme 2: Shared Experiences
Shared experiences—through active collaboration or simply being together—play a crucial role in forming meaningful relationships and lasting memories.
@@ 4.2.1. Fostering Shared Memories Through Collaborative Activities
Shared activities served as meaningful foundations for building friendships. P06 contrasted casual solo games with more immersive, co-experienced ones. While playing a simple emoji game lacked true shared experience since “we don’t actually see each other’s screens,” Roblox fostered lasting memories: “do you remember that time when that kid in the voice chat said this, or when we got scared by that jump scare… it’s not just that we’re in the same space, but that we’re doing something together and experiencing it together.”
P12 described the value of collaborative projects like “creat[ing] a house we love” online, which offered both “a sense of self-achievement” and “teamwork,” contrasting with the passivity of “just like doom scrolling.” Working together on “the same topic, on the same project” felt like “a better way to connect” than “posting” content individually.
@@ 4.2.2. Making Space for Casual Connections Through Low-Pressure, Ambient Social Interaction
Low-pressure, ambient interaction—co-existing in shared spaces or listening to music together—offered natural ways to build connections. P18 described a “study sesh” where friends could send avatars to “sit and study together” in a virtual room with Lo-Fi music, mirroring Discord servers’ 24/7 activity and offering comforting presence without direct conversation. Even “mindless scrolling” became meaningful when shared: “see[ing] their reaction, hav[ing] a discussion, and…commentat[ing] in real-time, instead of just sending it back and forth,” would feel “10 times more fulfilling.”
Music often facilitated these interactions. P19 envisioned a “jukebox” where friends could “queue up songs”:
@@
Music served as both “conversation starter” and passive connection: “you don’t have to connect through talking to someone about music. You can just listen to the music.”
4.3. Theme 3: Communal Ambience
SMH was envisioned as a platform for intentional, engaging interactions with close friends, making it naturally less prone to social comparison. P09 described it as “the antonym to pressure,” “stress-free,” “community oriented,” “wholesome,” “accessible,” and “accepting.” Unlike MSM, which often feels performative and stressful, SMH would foster organic, personalized interactions and closer friendships through neighborhood structures, shared gathering spaces, and stronger boundary controls.
@@ 4.3.1. Designing Neighborhood Structures that Reflect Real-World Communities
SMH’s 3D world would mirror real-world communities, using familiar layouts like homes, neighborhoods, and shared areas to foster interaction. P22 described this in detail:
@@
Physical distances within the landscape would represent social tie strength. One participant elaborated:
@@
P01 proposed a tiered layout where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be farther away, requiring a “little hike” to reach. This alignment of physical landscape with emotional closeness would streamline interactions. As P07 noted, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.”
@@ 4.3.2. Establishing Third Places for Casual Social Interaction and Serendipitous Connections
SMH neighborhoods would include shared third places where people could casually gather, lowering social barriers and creating connection opportunities. P05 described these as “neutral spaces” not tied to specific activities or relationships, with “some places to sit, and some kind of flat table-like surface” supporting casual activities like “doing homework or study, or [playing] some kind of board game.” Such spaces would function as “a third, neutral territory” for organic interaction without managing invitations or switching platforms.
Shared neighborhood spaces—common rooms, courtyards, clubhouses—would allow casual interaction outside personal spaces. P01 described gathering in a “courtyard or game room” at the neighborhood center. Drawing on Club Penguin (club-penguin) and Poptropica (FeigenbaumOtherPoptropicaGamesb), they imagined a “clubhouse… open to anybody,” with themed rooms catering to different interests, like a Quidditch field or library, creating opportunities to meet others through shared activities. P09 imagined a “virtual dining room” for casual conversation while eating—open and inclusive, enabling cross-social group connections unlike Hogwarts’ segmented house-based accommodation.
Public third places would also support serendipitous interactions. P03 described chance encounters while shopping: “you strike a conversation because you’re both picking up the same shirt… little connections like that happen just because you’re in a public space.” P21 imagined “small arenas” for “Quidditch games…[where users] can choose to get seated randomly [and] just start talking to the person next to you.”
@@ 4.3.3. Prioritizing Close Relationships and Reducing Extraneous Content
To preserve SMH’s focus on meaningful, stress-free interactions, participants emphasized filtering irrelevant content like celebrity drama or political commentary while prioritizing updates from close friends. P19 criticized how platforms like Instagram have shifted away from their initial focus:
@@
Several participants advocated minimizing celebrity influence, with P01 suggesting isolating them to a “celebrity zone.” P07 captured the preference: “If I’m just gonna open the app for like 5 seconds… I’m gonna wanna see my best friend’s post, not some girl I talked to once 3 years ago.” SMH would be “more designed for people you actually know… like your friends, classmates, or Quidditch team.”
@@ 4.3.4. Developing Socially Grounded and Themed Conversational Spaces
Themed, socially grounded conversational spaces for protected, meaningful dialogue could make SMH feel more personal than MSM.
@@
Many valued walking into themed, topic-specific rooms where conversation felt protected and purposeful, rather than broadcasting to broad, undefined audiences. P22 imagined “chat rooms you can walk into,” each centered on themes like “high school stuff,” decorated with artifacts from shared experiences. These would be “protected space[s]…where you can talk…without feeling unsafe or judged,” contrasting with MSM’s more exposed feel where “you don’t know everything around it.”
Personalizing these spaces would make them more emotionally connected. P07 described filling chatrooms with “photos of you together, inside jokes, or fun memories.” P09 imagined a “Room of Requirement”-like system (room-of-requirement) where users could create custom rooms on demand to match the “purpose” of their conversation. Interest-based spaces, similar to Discord (discord), would provide a “common point” (P13) or “good talking point” (P07) to break the ice.
@@ 4.3.5. Implementing Comprehensive Boundary Controls to Create Personal Spaces that Feel Safe
SMH would need comprehensive boundary controls beyond existing platforms’ limited and inconsistent privacy tools. P22 explained that even with contact syncing off, Instagram still suggested their Finsta (secondary account (VitakHuang-2022-FinstaAccounts-t)) to people they wanted to avoid, based on mutual follows. They stressed needing stronger blocking controls, including blocking “alternate accounts” and fully removing blocked person traces: “I don’t wanna be able to tell that the person I blocked is even around me.” P14 pointed out that even private Instagram accounts expose personal content like Story Highlights to “a random person”, revealing perceived privacy protection gaps (story-highlights).
Participants wanted stronger control over visibility, imagining features letting them opt out of being seen entirely. P03 suggested an “invisibility cloak” for moments when users “don’t want anyone to see you.”
@@
4.4. Theme 4: Nuanced Self-Presentation
Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14). This curated image culture creates “a lot of seriousness” (P23), reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23). P13 shared helping a friend prepare an Instagram post, including selecting “post order,” highlighting “how seriously we take social media and our image online.” SMH, by contrast, would allow self-presentation that feels more personal, authentic, and flexible—reflecting users’ full range without curation or performance pressures.
@@ 4.4.1. Expressing Identity Through Customizable Avatars, Spaces, and Artifacts
Customizable avatars and personal spaces offered meaningful self-expression far beyond MSM’s limited settings. P21 envisioned building their own “mini world” complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school,” sharing these spaces as “showing a little part of yourself.” P03 wanted to create a “dark academia” Hogwarts-inspired room, filled with “so many bookshelves and… Victorian architecture.”
Customizable “3D avatar[s]” (P11) and profiles were key ways of showing individuality. P22 reflected on early Myspace (myspace), describing how they enjoyed coding profile layouts and changing “the font and really make it reflect me.” Even small details like choosing usernames could “convey to the world and your friends your own identity… which is a positive feeling usually.” P01 noted such expression is often “mainly for you,” not just for public display. P17 added that even unconscious choices reveal something personal: “your brain conjured that up, right? So I think that’s part of you.”
Beyond visual customization, SMH would support ambient, indirect expression—subtle cues about what someone is doing, feeling, or thinking without formal posts. P01 wanted their avatar to reflect real-time states like current outfit or music, so friends could intuitively notice “‘Oh. that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ P10 imagined friends picking up on interest changes by observing virtual space updates over time, providing “an insight into… what I’m thinking.”
The contrast between 2D platforms’ flat feel and 3D environments’ immersive qualities mattered. P16 noted that VR creates presence feeling more like meeting “real people… compared to profile pictures,” because you can pick up on “mannerisms” and details like “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.”
These choices feel more authentic than generic posts because they emerge naturally from personal preferences. As P01 put it, such “creative expression” feels “a lot more personal” because “that’s not just like what life has given you… a reflection of your outlook on life.” P13 explained that these subtle forms of sharing make social media feel more like “an extension of everyday interaction,” helping people “gain the basic foremost information” about someone without direct conversation. P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
Personal spaces like bedrooms carry intimacy and authenticity that translates well to SMH. P13 explained:
@@
These ambient, low-pressure forms of expression also carried lower social stakes. P16, who described themselves as “fairly private” and not particularly “public” about their preferences, explained that “decorat[ing] a house” or customizing an avatar allows others to “get to know more about my personality… through my style” without feeling performative. This kind of expression felt “more justified” and less “embarrassing” than posting content just to be seen.